OpenStudy (goformit100):

The ratio between the length and the breadth of a rectangular park is 3 : 2. If a man cycling along the boundary of the park at the speed of 12 km/hr completes one round in 8 minutes, then the area of the park (in sq. m) is: A. 15360 B. 153600 C. 30720 D. 307200

OpenStudy (goformit100):

@.Sam.

OpenStudy (goformit100):

HELP sir

OpenStudy (goformit100):

@kropot72

OpenStudy (goformit100):

@LoveYou*69

OpenStudy (mayankdevnani):

let the length be 3x and breadth be 2x. ok @goformit100

OpenStudy (mayankdevnani):

so its perimeter would be \[=2(3x+2x)=10x\]

OpenStudy (mayankdevnani):

@kropot72 this answer would not match with the options!!!!

OpenStudy (mayankdevnani):

so its distance = \[\frac{12000}{60} \times 8=1600\]

OpenStudy (mayankdevnani):

so \[10x=1600\] \[x=160\] \[area=480 \times 320 =153600\]

OpenStudy (mayankdevnani):

so correct answer is option b @goformit100

OpenStudy (mayankdevnani):

m don't satisfy with @kropot72 answer!!!!!

OpenStudy (goformit100):

Thaks

OpenStudy (mayankdevnani):

welcome!!!!